### v1

@@ Appendix B Experimental setup
Figure M1: Wiring diagram. Experimental setup for control hardware, cable routing, and shielding for our device.
@@
Figure M2: Photo of device The device consists of a on-chip transmon and a co-planar waveguide resonator mounted inside a high-purity Aluminum cavity. The package is mounted to a gold-plated copper breadboard at the mixing-chamber plate of a dilution refridgerator.
@@
The device used in this paper consists of an oscillator, a 3D stub post cavity made from high-purity 4N Aluminum treated with an acid etch, and a transmon qubit. The transmon, made of Niobium, is fabricated on a resistive silicon chip, along with an on-chip readout resonator also made of Niobium. The single chip hosting the transmon and the readout resonator is mounted in the 3D cavity package using copper clamps. The cavity and the copper clamp contain copper films for thermalization directly to gold-plated copper breadboard at the mixing chamber plate of the dilution refrigerator (Fig. M2). The device is shielded with Copper coated with Berkeley Black, and two types of magnetic shields: Aluminum, and Cryoperm (Fig. M1).
@@
The control pulses for the qubit and the storage are synthetized using Zurich Instruments (ZI) HDAWG, which have a baseband bandwidth of 1 GHz. These are upconverted using Rohde & Schwarz SGS100A, which are signal generators with built-in IQ mixers. These built-in mixers are used for all frequency conversions with the exception of the readout. The readout pulses are synthesized and digitzied using ZI UHFQA, and are up-converted and down-converted using Marki mixers (MMIQ-0416LSM-2), with a split LO from a single SGS100A. Readout signals are first amplified with a traveling-wave Josephson Amplifier (TWPA), a quantum-limited amplifier. The TWPA typically requires large pump tones, so we gate it with a trigger line from the readout AWG which combines with the CW pump tone in an IQ mixer (as a makeshift fast switch). The readout signals are then further amplified with a High-electron mobility transistor (HEMT) ampliflier at the 4K stage, and again amplified with a room temperature amplifier (ZVA-1W-103+ from Mini-Circuits) and filtered. The digitizer on the ZI UHFQA converts to the analog response to a digital signal and integrates it to produce a binary outcome depending on the qubit state.
@@
Appendix C System Hamiltonian & Reservoir description
@@
We approximate our transmon as a qubit. Our qubit-oscillator system is well described by the Hamiltonian [Blais_2021]:
$H/\hbar=\omega_{q}q^{\dagger}q+\omega_{a}a^{\dagger}a-\chi q^{\dagger}qa^{\dagger}a-\chi^{\prime}q^{\dagger}qa^{\dagger 2}a^{2}-K_{q}q^{\dagger 2}q^{2}-Ka^{\dagger 2}a^{2}+\Omega(t)(q+q^{\dagger})+\epsilon(t)(a+a^{\dagger}),$ (18)
where $a$ is the annihilation operator for the oscillator mode, and $q$ is the annihilation operator for the qubit mode, $\omega_{a}$ and $\omega_{q}$ are the frequencies of the oscillator and qubit mode respectively, $\chi$ and $\chi^{\prime}$ are the dispersive shift and the cavity state-dependent dispersive shift respectively, $K$ and $K_{q}$ are the self-Kerr of the oscillator and the transmon anharmonicity respectively. The values for these parameters, as well as values for decay rates are listed in Table S1. For the construction of our drives, we ignore the self-Kerr or the oscillator as well as the higher-order cross-Kerr. We note that these are indeed present, but for the purposes of a quantum reservoir, only add to the complexity of the dynamics. Finally, moving to the rotating frame of the qubit and cavity mode, we arrive at the Hamiltonian in Eq. 1.
@@
Figure M3: Geometric phase unitary to sense phase of unknown displacement (a.) Decomposition of the unitary used throughout the reservoir for time-independent tasks. (b.) Schematic representation of the dynamics of the oscillator state under the reservoir drives with time-independent input, highlighting a unitary which implements a geometric phase unitary. (1) At the start of the protocol, the qubit is in the ground state, with the cavity at vacuum. (2) An initial $X_{\pi/2}$ pulse brings the qubit to the equator. The $\mathrm{CNOD}(\alpha)$ unitary conditions the state of the cavity based on the qubit. For the first reservoir, this is a coherent state. (3) and (4) For time independent inputs, the effective action can be described by a single displacement on the cavity mode. In this experiment, we operate the displacement at a frequency which, to first order, causes a displacement only on the state conditioned on the ground state of the qubit. A qubit $\pi$ pulses switches the state in between the two displacements. (5) The final conditional displacement brings the two conditioned states back onto each other. This effectively disentangles the qubit from the cavity mode. (6) The effective geometrical area enclosed $A$, which is a function of the input, is imparted onto the qubit.
@@
Whereas with a typical homodyne-setup, measuring the quadratures of some unknown signal is easy, however performing the same measurement of a displacement on an oscillator using only qubit measurements can be non-trivial. Of course, when designing a reservoir, one does not strive to implement the identity, but it is a good starting point – the unitary is thus implemented to approximate the identity. It consists of the input signal data, which is sandwiched on either side by fast conditional displacement gates implemented with CNOD [1] and qubit rotation gates. The broad-overview of the decomposed unitary is given in terms of gates in Fig. M3, along with a schematic portrayal of the phase-space trajectory of the oscillator mode initialized in vacuum subject to a time-independent drive.
@@
$U_{7}U_{6}U_{5}U_{4}U_{3}U_{2}=\frac{i}{\sqrt{2}}e^{\alpha\beta^{*}-\alpha^{*}\beta}D(\beta)(|g\rangle\langle g|-|e\rangle\langle g|)-\frac{i}{\sqrt{2}}e^{-\alpha\beta^{*}+\alpha^{*}\beta}D(\beta)(|g\rangle\langle e|+|e\rangle\langle e|)$ (26)
@@
where $A=2|\alpha||\beta|\sin(\delta)=i(\alpha\beta^{*}-\alpha^{*}\beta)$ is the geometric phase enclosed by the oscillator trajectory which is dependent on the phase difference $\delta$ between the known displacement $D(\alpha$), and the unknown displacement $D(\beta)$ (Fig M3b). Thus, for the proper qubit state before the application of $U_{2}\ldots U_{7}$, we are able to extract phase information of the displacement. We also note that the qubit and the oscillator are disentangled after the unitary, and that the effect of the unitary on the oscillator mode is a simple displacement.
@@
Thus, with this sequence of unitaries, we are able to extract the phase of some unknown displacement (relative to some known displacement $\alpha$) by simply measuring the qubit. While for the first run of the reservoir, the qubit will start in the ground state (up to thermal noise), after performing a parity measurement, the qubit state will depend on the previous measurement outcome. See Fig. M7 for an experimental implementation of the above results.
@@
Whereas all gates besides the input (Eqs. 21 and 23) are fast and therefore insensitive to the cross-Kerr interaction, the primary deviation from the gate description occurs for the input, which can be very long. This input displacement is conditioned on the qubit being in the ground state. Therefore, in the rotating from of the qubit-oscillator system, the branch of the cavity state conditioned on the qubit being in the excited state will rotate at a frequency $\chi$, which in general will break the geometric phase construction for time-independent tasks. Therefore, we limit the exposure time of the reservoir to the input signal to be an integer multiple of $4\pi/\chi$, so that the cavity state conditioned on the qubit being in the excited state will return to the same point.
@@
As mentioned above, the effect of the unitary on the oscillator state for time-independent signals is simply a displacement of the input data $D(\beta)$, independent of the qubit measurement outcome. For the following discussion, we will ignore the qubit dynamics, since the qubit and the oscillator are disentangled at the end of the unitary. In effect, the state of the cavity can be described by a series of alternating displacements and parity measurements:
@@
Figure M4: Schematic description of the dynamics of the oscillator, starting in vacuum, for a slow time-varying input In such a scenario (as is the case for the task of classifying radio modulation schemes), the signal causes a displacement largely conditioned on the ground state. Generally, the value of the displacement is different before and after the $\pi$ pulse in between (3) and (4). Unlike the regime for time-independent signals, there is no effective area enclosed in phase space, which leaves the qubit entangled with the cavity. (6) and (7) describe the state of the cavity after the qubit measurement. The resulting state of the cavity is a cat state, where the parity of the cat dependent on whether the outcome of the qubit measurement is ground or excited.
For generic, time-dependent signals, like those classified in Figs. 3 and 4 in the main text, the geometric unitary described by Eqs. 19-25 does not in general hold, as the symmetry between panels 3 and 4 in Fig. M3 is broken. Additionally, the approximation that the input is displacement conditioned on the qubit in the ground state (Eq. 21 and 23) will not hold for high bandwidth signals, like those in Fig. 4 in the main text. For high-bandwidth signals, the input will also have some contribution in displacing cavity state conditioned on the qubit being in the excited state, which can lead to complex dynamics in the cavity state. While for generic signals, this can be hard to describe, here we prove a treatment of our reservoir construction for slowly-varying, time-dependent signals, like those in Fig. 3 of the main text.
We can follow most of the derivation from the scenario of time independent signals in Appendix C, to describe the dynamics of the QRC for the task of classifying radio frequency modulation schemes. Along with the assumptions in the previous section, we make the slow-varying input approximation, such that the displacement on the oscillator of the reservoir is still effectively conditioned on the ground state of the qubit. The displacement on the cavity depends on the value of the symbol encoded for the given modulation scheme. Since, in general, the symbol is different before and after the qubit $\pi$ pulse: the direction of the displacement in the cavity will be different. Given the timescales of the input signal involved, this essentially corresponds to a displacement on the cavity state conditioned on the ground state of the qubit. When the two displacements are different in magnitude and direction, the qubit remains entangled with the cavity at the end of the reservoir unitary. The state of the system just before the measurements is (step (5) of Fig M4:
@@
where $\beta_{i}$ is the displacement before the qubit flip, and $\beta_{j}$ is the displacement after. $A_{i}=\alpha\beta_{i}$ is the phase acquired after two non-orthogonal displacements. When $\beta_{i}=\beta_{j}$, we recover the dynamics for time independent signals. It is straightforward to show that the qubit will be disentangled from the cavity and that the area $A_{i}$, corresponding to the geometrical phase form the area enclosed in phase space will be present as a relative phase difference between the ground and excited state. After a $Y_{\pi/2}$ gate, we have the following state in our system:
$|\psi\rangle=\frac{1}{2}[e^{iA_{i}}D(\beta_{i})+e^{-iA_{j}}D(\beta_{j})]\otimes|g,\mathrm{cavity}\rangle+\frac{1}{2}[e^{iA_{i}}D(\beta_{i})-e^{-iA_{j}}D(\beta_{j})]\otimes|e,\mathrm{cavity}\rangle,$ (35)
One can think of this as a cat state in the cavity, with a parity determined by the qubit state. This is schematically shown in (6) and (7) in Fig M4. In the limit of very different displacements, the probability of the qubit measurement is the same for both ground and excited states. The goal of this task can be thought of as discriminating probability distribution functions over the $(I,Q)$ plane. Fig 3 (a) represents the so-called “constellation ” diagram of the modulation schemes considered in this work. Each scheme can take discrete values in $(I,Q)$ space, with even equal probability (we construct the dataset of radio signals encoding random binary strings). Our lack of knowledge of the exact displacement on the cavity can be mathematically expressed as a density matrix. This is the most apparent in the state of the cavity after the initial qubit measurement,
$\rho_{\mathrm{cavity}}^{\prime}=\sum_{\beta_{i}\in P}p_{i}D^{\dagger}(\beta_{i})\rho_{\mathrm{cavity}}D(\beta_{i})+\sum_{\beta_{i},\beta_{j}\in P}p_{ij}e^{i(A_{i}-A_{j})}D^{\dagger}(\beta_{i})\rho_{\mathrm{cavity}}D(\beta_{j}),$ (36)
where $\rho_{\mathrm{cavity}}^{\prime}$ is the density matrix representation of the cavity right after the qubit measurement and $\rho_{\mathrm{cavity}}$ describes the initial density matrix before the application of the input. The set $P$ describes the distribution of possible displacements which can be received from the input. $p_{i}$ is probability for receiving the symbol corresponding to a displacement $\beta_{i}$, and $p_{ij}$ is the conditional probability of displacement $\beta_{j}$, given $\beta_{i}$. For the task considered in this work, these distributions are uniform, with no contributions from conditional probabilities. However, this description of the reservoir motivates the potential for the QRC to distinguish signals with complex correlations in the symbols of the message encoded.
Appendix D Quantum reservoir characterization

### v2

@@ Appendix A Experimental setup
Figure S1: Wiring diagram. Experimental setup for control hardware, cable routing, and shielding for our device.
@@
Figure S2: Photo of device The device consists of a on-chip transmon and a co-planar waveguide readout resonator mounted inside a high-purity Aluminum cavity. The package is mounted to a gold-plated copper breadboard at the mixing-chamber plate of a dilution refrigerator.
@@
The device used in this paper consists of an oscillator, a 3D stub post cavity made from high-purity 4N Aluminum treated with an acid etch, and a transmon qubit. The transmon, made of Niobium, is fabricated on a resistive silicon chip, along with an on-chip readout resonator also made of Niobium. The single chip hosting the transmon and the readout resonator is mounted in the 3D cavity package using copper clamps. The cavity and the copper clamp contain copper films for thermalization directly to gold-plated copper breadboard at the mixing chamber plate of the dilution refrigerator (Fig. S2). The device is shielded with Copper coated with Berkeley Black, and two types of magnetic shields: Aluminum, and Cryoperm (Fig. S1). The cavity pin is set such that the oscillator mode is undercoupled to the transmission line by a factor of 40. While this reduces the transmission of photons incident on our device by a factor of 40, it keeps the oscillator state thermalized to the fridge rather than the transmission line.
@@
The control pulses for the qubit and the storage are synthesized using Zurich Instruments (ZI) HDAWG, which have a baseband bandwidth of 1 GHz. These are upconverted using Rohde & Schwarz SGS100A, which are signal generators with built-in IQ mixers. These built-in mixers are used for all frequency conversions with the exception of the readout. The readout pulses are synthesized and digitized using ZI UHFQA, and are up-converted and down-converted using Marki mixers (MMIQ-0416LSM-2), with a split LO from a single SGS100A. Readout signals are first amplified with a traveling-wave Josephson Amplifier (TWPA), which is a quantum-limited amplifier. The TWPA typically requires large pump tones, so we gate it with a trigger line from the readout AWG which combines with the CW pump tone in an IQ mixer (as a makeshift fast switch). The readout signals are then further amplified with a High-electron mobility transistor (HEMT) ampliflier at the 4K stage, and again amplified with a room temperature amplifier (ZVA-1W-103+ from Mini-Circuits) and filtered. The digitizer on the ZI UHFQA converts to the analog response to a digital signal and integrates it to produce a binary outcome depending on the qubit state.
@@
Appendix B System Hamiltonian & Reservoir description
@@
Our transmon-cavity system is well approximated by the Hamiltonian [Blais_2021]:
$H/\hbar=\omega_{q}q^{\dagger}q+\omega_{a}a^{\dagger}a-\chi q^{\dagger}qa^{\dagger}a-\chi^{\prime}q^{\dagger}qa^{\dagger 2}a^{2}-K_{q}q^{\dagger 2}q^{2}-Ka^{\dagger 2}a^{2}+{{\Xi}}(t)(q+q^{\dagger})+{{\xi}}(t)(a+a^{\dagger}),$ (18)
where $a$ is the annihilation operator for the oscillator mode, and $q$ is the annihilation operator for the qubit mode, $\omega_{a}$ and $\omega_{q}$ are the frequencies of the oscillator and qubit mode respectively, $\chi$ and $\chi^{\prime}$ are the dispersive shift and the oscillator state-dependent dispersive shift respectively, $K$ and $K_{q}$ are the self-Kerr of the oscillator and the transmon anharmonicity respectively. The values for these parameters, as well as values for decay rates, are listed in Table S1. The last two terms describe the qubit and oscillator drives in the lab frame. The lab-frame drives are related to the rotating-frame drives in Eq. 1 via $\Xi(t)=\Omega(t)e^{i\omega_{q}t}+\mathrm{H.c.}$ and $\xi(t)=\epsilon(t)e^{i\omega_{a}t}+\mathrm{H.c.}$. For the design of our drives, we ignore the self-Kerr of the oscillator as well as the higher-order cross-Kerr. We note that these are indeed present, but for the purposes of a quantum reservoir, only add to the complexity of the dynamics. Finally, moving to the rotating frame of the transmon and oscillator mode and truncating to the first two levels of the transmon, we arrive at the Hamiltonian in Eq. 1.
@@
Figure S3: Geometric phase unitary to sense phase of unknown displacement (a.) Decomposition of the unitary used throughout the reservoir for time-independent tasks. (b.) Schematic representation of the dynamics of the oscillator state under the reservoir drives with time-independent input, highlighting a unitary which implements a geometric phase unitary. (1) At the start of the protocol, the qubit is in the ground state, with the oscillator at vacuum. (2) An initial $X_{\pi/2}$ pulse brings the qubit to the equator. The $\mathrm{CNOD}(\alpha)$ unitary conditions the state of the oscillator based on the qubit. For the first reservoir, this is a coherent state. (3) and (4) For time independent inputs, the effective action can be described by a single displacement on the oscillator mode. In this experiment, we operate the displacement at a frequency which, to first order, causes a displacement only on the state conditioned on the ground state of the qubit. A qubit $\pi$ pulses switches the state in between the two displacements. (5) The final conditional displacement brings the two conditioned states back onto each other. This effectively disentangles the qubit from the oscillator mode. (6) The effective geometrical area enclosed $A$, which is a function of the input, is imparted onto the qubit.
@@
While measuring the quadratures of some unknown signal is easy with a typical homodyne setup, performing the same measurement of a displacement on an oscillator using only qubit measurements can be non-trivial. Of course, when designing a reservoir, one does not strive to implement the identity, but it is a good starting point – the unitary is thus implemented to approximate the identity. It consists of the input signal data, which is sandwiched on either side by fast conditional displacement gates implemented with CNOD [1] and qubit rotation gates. The broad-overview of the decomposed unitary is given in terms of gates in Fig. S3, along with a schematic portrayal of the phase-space trajectory of the oscillator mode initialized in vacuum subject to a time-independent drive.
@@
$U_{7}U_{6}U_{5}U_{4}U_{3}U_{2}=\frac{i}{\sqrt{2}}e^{(\alpha\beta^{*}-\alpha^{*}\beta)}D(\beta)(|g\rangle\langle g|-|e\rangle\langle g|)-\frac{i}{\sqrt{2}}e^{(-\alpha\beta^{*}+\alpha^{*}\beta)}D(\beta)(|g\rangle\langle e|+|e\rangle\langle e|)$ (26)
@@
where $A=2|\alpha||\beta|\sin(\delta)=i(\alpha\beta^{*}-\alpha^{*}\beta)$ is the geometric phase enclosed by the oscillator trajectory which is dependent on the phase difference $\delta$ between the known displacement $D(\alpha$), and the unknown displacement $D(\beta)$ (Fig S3b). Thus, for the proper qubit state before the application of $U_{2}\ldots U_{7}$, we are able to extract information about the phase of the displacement. We also note that the qubit and the oscillator are disentangled after the unitary, and that the effect of the unitary on the oscillator mode is a simple displacement.
@@
Thus, with this sequence of unitaries, we are able to extract the phase of some unknown displacement (relative to some known displacement $\alpha$) by simply measuring the qubit. While for the first run of the reservoir, the qubit will start in the ground state (up to thermal noise), after performing a parity measurement, the qubit state will depend on the previous measurement outcome. See Fig. S7 for an experimental implementation of the above results.
@@
Whereas all gates besides the input (Eqs. 21 and 23) are fast and therefore insensitive to the cross-Kerr interaction, the primary deviation from the gate description occurs for the input, which can be very long. This input displacement is conditioned on the qubit being in the ground state. Therefore, in the rotating from of the qubit-oscillator system, the branch of the oscillator state conditioned on the qubit being in the excited state will rotate at a frequency $\chi$, which in general will break the geometric phase construction that works for time-independent tasks. Therefore, we limit the exposure time of the reservoir to the input signal to be an integer multiple of $4\pi/\chi$, so that the oscillator state conditioned on the qubit being in the excited state will return to the same point.
@@
As mentioned above, the effect of the unitary on the oscillator state for time-independent signals is simply a displacement of the input data $D(\beta)$, independent of the qubit measurement outcome. For the following discussion, we will ignore the qubit dynamics, since the qubit and the oscillator are disentangled at the end of the unitary. In effect, the state of the oscillator can be described by a series of alternating displacements and parity measurements:
@@
Figure S4: Schematic description of the dynamics of the oscillator, starting in vacuum, for a slow time-varying input In such a scenario (as is the case for the task of classifying radio modulation schemes), the signal causes a displacement largely conditioned on the ground state. Generally, the value of the displacement is different before and after the $\pi$ pulse in between (3) and (4). Unlike the regime for time-independent signals, there is no effective area enclosed in phase space, which leaves the qubit entangled with the oscillator. (6) and (7) describe the state of the oscillator after the qubit measurement. The resulting state of the oscillator is a cat state, where the parity of the cat dependent on whether the outcome of the qubit measurement is ground or excited.
For generic, time-dependent signals, like those classified in Figs. 3 and 4 in the main text, the geometric unitary described by Eqs. 19-25 does not in general hold, as the symmetry between panels 3 and 4 in Fig. S3 is broken. Additionally, the approximation that the input is displacement conditioned on the qubit in the ground state (Eq. 21 and 23) will not hold for high bandwidth signals, like those in Fig. 4 in the main text. For high-bandwidth signals, the input will also have some contribution in displacing oscillator conditioned on the qubit being in the excited state, which can lead to complex dynamics in the oscillator. While for generic signals, this can be hard to describe, here we prove a treatment of our reservoir construction for slowly-varying, time-dependent signals, like those in Fig. 3 of the main text.
We can follow most of the derivation from the scenario of time independent signals in Appendix B, to describe the dynamics of the QRC for the task of classifying radio frequency modulation schemes. Along with the assumptions in the previous section, we make the slow-varying input approximation, such that the displacement on the oscillator of the reservoir is still effectively conditioned on the ground state of the qubit. The displacement on the oscillator depends on the value of the symbol encoded for the given modulation scheme. Since, in general, the symbol is different before and after the qubit $\pi$ pulse: the direction of the displacement in the oscillator will be different. Given the timescales of the input signal involved, this essentially corresponds to a displacement on the oscillator conditioned on the ground state of the qubit. When the two displacements are different in magnitude and direction, the qubit remains entangled with the oscillator at the end of the reservoir unitary. The state of the system just before the measurements is (step (5) of Fig S4:
@@
where $\beta_{i}$ is the displacement before the qubit flip, and $\beta_{j}$ is the displacement after. $A_{i}=2\text{Im}\left[\alpha\beta_{i}^{*}\right]$ is the phase acquired after two non-orthogonal displacements. When $\beta_{i}=\beta_{j}$, we recover the dynamics for time independent signals. It is straightforward to show that the qubit will be disentangled from the oscillator and that the area $A_{i}$, corresponding to the geometrical phase form the area enclosed in phase space will be present as a relative phase difference between the ground and excited state. After a $Y_{\pi/2}$ gate, we have the following state in our system:
$|\psi\rangle=\frac{1}{2}[e^{iA_{i}}D(\beta_{i})+e^{-iA_{j}}D(\beta_{j})]|g,\mathrm{cavity}\rangle+\frac{1}{2}[e^{iA_{i}}D(\beta_{i})-e^{-iA_{j}}D(\beta_{j})]|e,\mathrm{cavity}\rangle,$ (35)
One can think of this as a cat state in the cavity, with a parity determined by the qubit state. This is schematically shown in (6) and (7) in Fig S4. In the limit of very different displacements, the probability of the qubit measurement is the same for both ground and excited states. The goal of this task can be thought of as discriminating probability distribution functions over the $(I,Q)$ plane. Fig 3 (a) represents the so-called “constellation ” diagram of the modulation schemes considered in this work. Each scheme can take discrete values in $(I,Q)$ space, with even equal probability (we construct the dataset of radio signals encoding random binary strings). Our lack of knowledge of the exact displacement on the oscillator can be mathematically expressed as a density matrix. This is the most apparent in the state of the oscillator after the initial qubit measurement,
$\rho_{\mathrm{cavity}}^{\prime}=\sum_{\beta_{i}\in P}p_{i}D(\beta_{i})\rho_{\mathrm{cavity}}^{\dagger}D(\beta_{i}),$ (36)
where $\rho_{\mathrm{cavity}}^{\prime}$ is the density matrix representation of the cavity right after the qubit measurement and $\rho_{\mathrm{cavity}}$ describes the initial density matrix before the application of the protocol. The set $P$ describes the distribution of possible displacements which can be received from the input. $p_{i}$ is probability for receiving the symbol corresponding to a displacement $\beta_{i}$ For the task considered in this work, these distributions are uniform, with no contributions from conditional probabilities. However, this description of the reservoir motivates the potential for the QRC to distinguish signals with complex correlations in the symbols of the message encoded.
Appendix C Quantum reservoir characterization